OpenStudy (anonymous):

What percent of 80 is 24?

OpenStudy (anonymous):

24/80 should get you your answer. If I had $24 dollars, that would be 24% of $100, right? By dividing 24/100, I should get that percent. So just replace 100 with 80.

OpenStudy (anonymous):

Let the percentage be x. So, (x/100)*80=24 x/100=24/80 Divide 24/80 by 8, we have: x/100=3/10 10x=300 Or, x=300/10 So, x=30 is your answer :)